OpenStudy (anonymous):

Factor each trinomial below. Please show your work. 1.x2 – 8x + 15 2.a2 – a – 20 3.a2 – 5a – 20 4.a2 + 12ab + 27b2 5.2a2 + 30a + 100

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started

OpenStudy (anonymous):

ok thanks!

jimthompson5910 (jim_thompson5910):

you need to find two numbers that multiply to 15 (last term) and add to -8 (middle coefficient) these two numbers are -3 and -5 since -3*(-5) = 15 and -3 + (-5) = -8 so this means x^2 - 8x + 15 factors to (x-3)(x-5)

jimthompson5910 (jim_thompson5910):

this trick only works if the leading coefficient is 1

OpenStudy (anonymous):

ok so it will work for the first 4?

jimthompson5910 (jim_thompson5910):

correct, #4 is a bit different but it's basically the same as the first 3

OpenStudy (anonymous):

ok now what about the last one?

jimthompson5910 (jim_thompson5910):

for #4, instead of saying (a+NUMBER1)(a+NUMBER2) you would say (a+NUMBER1*b)(a+NUMBER2*b) where NUMBER1 and NUMBER2 are the two numbers you found

OpenStudy (anonymous):

then what would 'b' be?

OpenStudy (anonymous):

the third number?

jimthompson5910 (jim_thompson5910):

because 2 goes into each coefficient, you can factor out 2 to get 2a^2 + 30a + 100 2(a^2 + 15a + 50) now factor a^2 + 15a + 50 and keep the 2 outside in mind

jimthompson5910 (jim_thompson5910):

no b is just a variable

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

look at #4 and notice you have the two variables a and b

OpenStudy (anonymous):

ya but there is no sqr. rt of 50 but you could /2 again...?

OpenStudy (anonymous):

*#5

OpenStudy (anonymous):

ok...

jimthompson5910 (jim_thompson5910):

2a^2 + 30a + 100 2(a^2 + 15a + 50) 2(a+10)(a+5) ... since 10+5 = 15 and 10*5 = 50

jimthompson5910 (jim_thompson5910):

see how I'm getting that?

OpenStudy (anonymous):

oh ok so that would be the answer right?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

ok Now that makes sense! Thanks!

jimthompson5910 (jim_thompson5910):

yw